A population increased to P0e4 in the first four years. What is the growth rate of this population? [Hint: P(t) = P0ekt]
Answer by fractalier(6550)

From P(t) = P(0)e^kt, we have
P(4) = P(0)e^4k = P(0)e^4
Thus k must be 1.